The actuary for the pension plan of Carla Inc. calculated the following net gains and losses. Incurred during the Year (Gain)Year Minimum Amortization of (Gain) Loss | 2020 0 2021 0 2022 0 2023 0
Answer #1

First, we calculate the Average remaining service life per employee.

Average remaining service life per employee = Expected future years of service/Number of employees

Average remaining service life per employee = 4800/400

Average remaining service life per employee = 12

Next, we find the amortization of unrecognized net gain or loss

(Gain) or Loss For the Year Ended December 31,

Amount

2020

$302100

2021 $479800
2022 -$208400
2023 -$288100

Year

Projected benefit obligation (a)

Plan Assets (a)

Corridor (b)

Cumulative Unrecognized (Gain) Loss (a)

Minimum Amortization of (Gain) Loss

2020 $4006100 $2395800 $400610

0

0

2021

$4486000

2,204,800

$448600

302100

0

2022

$4971200

2,601,700

$497120

781900

23732

2023

$4238300

3,046,400

$423830

549768

10495

Explanations:

(a) As of the beginning of the year.

(b) The corridor is 10 percent of the greater of projected benefit obligation or plan assets.

(c) $781900– $497120 = $284780; $284780/12 = $23732

(d) $781900 – $23732 – $208400 = $549768

(e)  $549768 – $423830 = $125938; $125938/12 = $10495